OpenStudy (anonymous):

x = c tan t and y = c cot t find dy/dx ( ans given : - c^2/x^2 , I got - cot ^2 (t) or - y^2/c^2 , which is correct?)

ganeshie8 (ganeshie8):

both are right

ganeshie8 (ganeshie8):

since, (tant) * (cot t) = 1 that gives : (x/c)(y/c) = 1 xy = c^2

ganeshie8 (ganeshie8):

u can jump from x to y or vice versa using above relation !

OpenStudy (anonymous):

thanks @ganeshie8